# Lee-Yang theory of the superradiant phase transition in the open Dicke model

### v1

@@ I Introduction
Figure 1: Lee-Yang theory of the open Dicke model. (a) The system consists of an optical cavity with resonance frequency $\omega_{c}$ that is coupled to $N$ two-level atoms with energy splitting $\hbar\omega_{a}$. The coupling amplitude is denoted by $\lambda$, while $\kappa$ is the rate at which photons are emitted from the cavity. (b) Fourth (dashed, $k=4$) and fifth (solid, $k=5$) factorial cumulant of the photon emission statistics in the normal phase. (c) Poles of the factorial cumulant generating function extracted from the results in panel (b). The black dots mark the convergence points in the long-time limit. (d) Extracted convergence points as a function of the coupling. Parameters are $\omega_{c}=2\kappa$ and $\omega_{a}=0.5\kappa$, and $\lambda_{0}$ denotes the critical point of the closed Dicke model, while $\lambda_{-}$ and $\lambda_{+}$ are the critical points of the open Dicke model according to Eqs. (12, 22).
To observe the superradiant phase transition, effective physical realizations involving driven-dissipative systems have been explored. In these setups, internal [20, 21] or motional [22, 23] degrees of freedom are pumped to realize effective strong couplings, but with finite dissipation rates. The combination of the external driving and dissipation results in a non-equilibrium phase transition, which is of a different universality class than for the closed system [20, 24, 4]. However, this phase transition is easier to access experimentally, and it will be the focus of our work. The phase transition occurs in the thermodynamic limit of many atoms and is visible in the photon emission statistics collected over a long time. For example, the cavity occupation changes abruptly at the critical coupling [25]. At the same, however, measurements of the photon counting statistics are of a finite duration, so practical approaches to predict the long-time behavior from finite-time statistics are called for.
@@ II Dicke model
Figure 1(a) shows a physical implementation of the Dicke model consisting of a single-mode optical cavity coupled to an ensemble of two-level atoms. Other realizations include superconducting circuits [5, 6, 7, 8, 9, 10, 11, 12, 13, 14], trapped ions [15], and collective electronic systems [16]. Some implementations have been based on driving internal transitions of multi-level atoms [20, 21], or on the motional states of atoms in a Bose-Einstein condensate [22, 23]. The latter implementation involves dissipative effects, which we will also include in our treatment.
@@ IV Lee-Yang theory
While the theory of Lee and Yang was developed for equilibrium phase transitions [26, 27], it has subsequently been extended to several types of non-equilibrium situations [30, 28, 29]. Those include dynamical phase transitions (also known as trajectory phase transitions) [31, 41], dynamical quantum phase transitions in quantum many-body systems after a quench [32, 33, 34, 35], as well as to quantum phase transitions at zero temperature [36, 37, 38]. In such nonequilibrium situations, the phase transitions are not manifested in the partition function or the associated free-energy density. Instead, other quantities, such as the moment generating function or the Loschmidt amplitude, play the role of the partition function, whose complex zeros determine the phase behavior of the system.
In the approach that we follow here, the generating function in Eq. (23) plays the role of the partition function. The partition function is given by a sum over all possible microconfigurations of an equilibrium system weighted by Bolzmann factors. Similarly, the factorial moment generating function is given by a sum over all possible dynamical trajectories, characterized by the number $n=0,1,2\ldots$ of photons that have been emitted during the time span $[0,t]$, weighted by the counting variable $s$. For an equilibrium system, the thermodynamic limit is approached as the system size, for example, the number of spins in a spin lattice, is increased. By contrast, in our case, the thermodynamic limit is approached as the observation time is increased, and the quantum jump trajectories become long. Thus, as our dynamical free-energy density, we take the scaled factorial cumulant generating function, defined in the long-time limit as
@@ V Phase transitions
We now extract the pole that is closest to the origin of the complex plane of the counting variable from the high factorial cumulants. This procedure is illustrated in Fig. 2(a), where we show the fourth and fifth factorial cumulants as a function of the observation time for three different couplings in the normal phase. For each coupling, we extract the position of the dominant pole, which we show in Fig. 2(b) as a function of the inverse observation time. The position of the dominant pole is well-approximated by the expression
$|s_{0}(t)-s_{c}|=1/\kappa t,$ (47)
where $s_{c}$ is the convergence point in the limit of long times. Thus, for each value of the coupling, we can extract the convergence point in the long-time limit, and in Fig. 2(c), we show the extracted convergence point as a function of the coupling. For small couplings, the convergence point does not reach the origin. However, as we approach the critical coupling, $\lambda\simeq\lambda_{-}$, the convergence point becomes smaller, and it eventually vanishes at $\lambda=\lambda_{-}$, signaling a phase transition. Thus, from measurements of the forth and fifth factorial cumulants at finite times, one can detect the phase transition at $\lambda=\lambda_{-}$. In particular, from Figs. 2(a,b), we see that one would be able to determine the convergence point in the long-time limit from measurements of the factorial cumulants for times that are shorter than $\kappa t=1$.
@@ VI Conclusions and outlook
We have seen that the system transits directly from the normal phase to the superradiant phase at a specific critical coupling, if the emission rate from the cavity is small. However, for larger decay rates, the critical coupling splits into two different values, and a small gap develops between the two phases. Here, we have applied our Lee-Yang theory to each of the phases, where the system can be mapped onto two uncoupled quantum harmonic oscillators, and the problem can be solved. By contrast, we are not aware of a solution in the gap region, which would allow us to calculate the factorial cumulants and thereby apply our Lee-Yang method. However, we expect that it may be possible to explore this region using a different diagonalization procedure [43], higher-order operator-cumulant expansions beyond the mean-field approximation [4], or advanced numerical methods for systems of finite size. A combination of such approaches might make it possible to explore the region between the two phases using our Lee-Yang formalism. We leave this task as an interesting open problem for future work.

### v2

@@ I Introduction
Figure 1: Lee-Yang theory of the open Dicke model. (a) The system consists of an optical cavity with resonance frequency $\omega_{c}$ that is coupled to $N$ two-level atoms with energy splitting $\hbar\omega_{a}$. The coupling amplitude is denoted by $\lambda$, while $\kappa$ is the rate at which photons are emitted from the cavity. (b) Fourth (dashed, $k=4$) and fifth (solid, $k=5$) factorial cumulant of the photon emission statistics in the normal phase defined in Eqs. (23-27). (c) Poles of the factorial cumulant generating function extracted from the results in panel (b) using Eq. (46). The black dots mark the convergence points in the long-time limit, $s_{c}$. (d) Extracted convergence points as a function of the coupling. Parameters are $\omega_{c}=2\kappa$ and $\omega_{a}=0.5\kappa$, and $\lambda_{0}$ denotes the critical point of the closed Dicke model, while $\lambda_{-}$ and $\lambda_{+}$ are the critical points of the open Dicke model according to Eqs. (12, 22).
To observe the superradiant phase transition, effective physical realizations involving driven-dissipative systems have been explored. In these setups, internal [20, 21] or motional [22, 23] degrees of freedom are pumped to realize effective strong couplings, but with finite dissipation rates. The combination of the external driving and dissipation results in a non-equilibrium phase transition, which is of a different universality class than for the closed system [20, 24, 4, 25, 26]. However, this phase transition is easier to access experimentally, and it will be the focus of our work. The phase transition occurs in the thermodynamic limit of many atoms and is visible in the photon emission statistics collected over a long time. For example, the cavity occupation changes abruptly at the critical coupling [27]. At the same time, however, measurements of the photon counting statistics are of a finite duration, so practical approaches to predict the long-time behavior from finite-time statistics are needed.
@@ II Dicke model
Figure 1(a) shows a physical implementation of the Dicke model consisting of a single-mode optical cavity coupled to an ensemble of two-level atoms, noting that other physical realizations are also possible [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. Some implementations are based on driving internal transitions of multi-level atoms [20, 21], or on the motional states of atoms in a Bose-Einstein condensate [22, 23]. The latter implementation involves dissipative effects, which we will also include in our treatment.
@@ IV Lee-Yang theory
While the theory of Lee and Yang was developed for equilibrium phase transitions [28, 29], it has subsequently been extended to several types of non-equilibrium situations [32, 30, 31, 33, 34, 35, 36, 37, 38, 39, 40]. In such nonequilibrium situations, the phase transitions are not manifested in the partition function or the associated free-energy density. Instead, other quantities, such as the moment generating function or the Loschmidt amplitude, play the role of the partition function, whose complex zeros determine the phase behavior of the system. In the approach that we follow here, the generating function in Eq. (23) plays the role of the partition function. The partition function is given by a sum over all possible microconfigurations of an equilibrium system weighted by Boltzmann factors. Similarly, the factorial moment generating function is given by a sum over all possible dynamical trajectories, characterized by the number $n=0,1,2\ldots$ of photons that have been emitted during the time span $[0,t]$, weighted by the counting variable $s$. For an equilibrium system, the thermodynamic limit is approached as the system size, for example, the number of spins in a spin lattice, is increased. By contrast, in our case, the thermodynamic limit is approached as the observation time is increased, and the quantum jump trajectories become long. Thus, as our dynamical free-energy density, we take the scaled factorial cumulant generating function, defined in the long-time limit as
@@ V Phase transitions
We now extract the pole that is closest to the origin of the complex plane of the counting variable from the factorial cumulants. This procedure is illustrated in Fig. 2(a), where we show the fourth and fifth factorial cumulants as a function of the observation time for three different couplings in the normal phase. We have checked that the results do not significantly change, if the cumulant order is increased, which shows that the approximations in Eqs. (45,46) are valid. For each coupling, we extract the position of the dominant pole, which we show in Fig. 2(b) as a function of the inverse observation time. The position of the pole is well-approximated by the expression [43, 55, 59, 56]
$|s_{0}(t)-s_{c}|\propto 1/\kappa t,$ (47)
where $s_{c}$ is the convergence point in the limit of long times. Thus, for each value of the coupling, we can extract the convergence point in the long-time limit, and in Fig. 2(c), we show the extracted convergence point as a function of the coupling. For small couplings, the convergence point does not reach the origin. However, as we approach the critical coupling, $\lambda\simeq\lambda_{-}$, the convergence point becomes smaller, and it eventually vanishes at $\lambda=\lambda_{-}$, signaling a phase transition. Thus, from measurements of the forth and fifth factorial cumulants at finite times, one can detect the phase transition at $\lambda=\lambda_{-}$. In particular, from Figs. 2(a,b), we see that one would be able to determine the convergence point in the long-time limit from measurements of the factorial cumulants for times that are shorter than $\kappa t=1$. Here, it should be noted that the extraction of the dominant pole in Eq. (46) works best for short times, where the poles are well-separated. Also, high factorial cumulants can be more accurately measured at short times [43, 52].
@@ VI Conclusions and outlook
We have seen that the system transits directly from the normal phase to the superradiant phase at a specific critical coupling, if the emission rate from the cavity is small. However, for larger decay rates, the critical coupling splits into two different values, and a small gap develops between the two phases. Here, we have applied our Lee-Yang theory to each of the phases, where the system can be mapped onto two uncoupled quantum harmonic oscillators, and the problem can be solved. By contrast, we are not aware of a solution in the gap region, which would allow us to calculate the factorial cumulants and thereby apply our Lee-Yang method. However, theoretically, it may be possible to explore this region using a different diagonalization procedure [45], higher-order operator-cumulant expansions beyond the mean-field approximation [4], or advanced numerical methods for systems of finite size. Also, measurements of the photon counting statistics for an open Dicke model may shed further light on the unexplored region between the two critical points using our Lee-Yang formalism. We leave these tasks as interesting open problems for future work.